OpenStudy (anonymous):

plz help me i have been asking since yesterday -6/7y-6≥42

OpenStudy (anonymous):

\[\frac{ -6 }{ 7 }y-6 \leq 42\]

OpenStudy (anonymous):

ok than wat?

OpenStudy (anonymous):

thats the question

OpenStudy (anonymous):

add 6 both sides

OpenStudy (anonymous):

ok than wat

OpenStudy (anonymous):

\[-48 \leq \frac{ 6 }{ 7 }y\]@Hero is this how you would ignore -multiplication

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):